### v5

# TranSUN: A Preemptive Paradigm to Eradicate Retransformation Bias Intrinsically from Regression Models in Recommender Systems

Abstract
Regression models are crucial in recommender systems. However, retransformation bias problem has been conspicuously neglected within the community. While many works in other fields have devised effective bias correction methods, all of them are post-hoc cures externally to the model, facing practical challenges when applied to real-world recommender systems. Hence, we propose a preemptive paradigm to eradicate the bias intrinsically from the models via minor model refinement. Specifically, a novel TranSUN method is proposed with a joint bias learning manner to offer theoretically guaranteed unbiasedness under empirical superior convergence. It is further generalized into a novel generic regression model family, termed Generalized TranSUN (GTS), which not only offers more theoretical insights but also serves as a generic framework for flexibly developing various bias-free models. Comprehensive experimental results demonstrate the superiority of our methods across data from various domains, which have been successfully deployed in two real-world industrial recommendation scenarios, i.e. product and short video recommendation scenarios in Guess What You Like business domain in the homepage of Taobao App (a leading e-commerce platform with DAU > 300M), to serve the major online traffic.
1 Introduction
Regression models play a pivotal role in recommender systems, which have been widely applied to prediction tasks for targets including user engagement duration [38, 9, 23, 22], transaction amount [15, 42, 44], lifetime value [26, 41, 40, 21, 24], etc. Mean squared error (MSE) loss is a predominant choice for loss functions of regression models, which adopts for target $y$ and features $x$ an apriori assumption of $y|x\sim\mathcal{N}(f(x;\theta),\sigma^{2})$, where $\mathcal{N}(f(x;\theta),\sigma^{2})$ is a Gaussian distribution with $f(x;\theta)$ as mean and $\sigma^{2}$ as variance and $f(x;\theta)$ is a $\theta$-parameterized predictor aiming to learn $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]$. However, the real target distribution frequently violates the Gaussian assumption in recommendation scenarios due to the high-skewed characteristics of the target data [7, 41, 40], which critically impedes the model convergence. This problem motivates prevalent adoption of various transformations $\mathrm{T}$ for target $y$ in the MSE-based regression models, abbr. transformed MSE, which facilitates the conformity between the transformed distribution $\mathcal{P}\big(\mathrm{T}(\mathrm{Y})|x\big)$ and the hypothetical distribution $\mathcal{N}(f(x;\theta),\sigma^{2})$, thereby improving the performance of model convergence [32, 31].
Though effective, it leads to the retransformation bias problem [2, 31, 37], i.e. the estimate for $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]$ is biased when the prediction $f(x;\theta)$ of the transformed MSE model is back-transformed to the original space, as
$\mathrm{T}^{-1}\big(f(x;\theta)\big)\neq\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y],$ (1)
where $\mathrm{T}^{-1}$ is the inverse function of $\mathrm{T}$. This critical issue has been conspicuously neglected within the community of recommender systems. Fortunately, there are still many effective works in other
Figure 1: Illustration of biasedness and convergence of MSE, LogMSE, and our TranSUN ($\mathrm{T}$ is logarithmic) during the training on the Indus dataset, where PGR denotes the ratio of the prediction mean to the ground truth mean in a batch and the loss value is normalized by the maximum value. As illustrated, MSE model has difficulty converging, while LogMSE model is significantly biased (i.e. PGR < 1.0). Differently, our method simultaneously maintains prediction unbiasedness (i.e. PGR is near 1.0) and superior convergence.
fields specifically studying the bias problem, all of which resort to the statistics of the model residuals to design post-processing bias correction mechanisms exactly after the model has been trained [12, 13, 33, 31, 14]. Despite the progress, all the prior works are post-hoc approaches designed and applied externally to the model, which face practical challenges when applied to the real-world recommender systems (detailed in Appendix C). Therefore, we challenge the prior paradigm by asking: Can we eradicate the retransformation bias intrinsically from the model in a preemptive manner?
To answer this, this paper proposes a novel preemptive paradigm to eradicate bias from the model exactly during the training phase via minor model refinement. Specifically, we propose the novel TranSUN method to empower the transformed MSE model with intrinsic unbiasedness via introducing a jointly-learned auxiliary model branch for explicit bias modeling. Our method not only achieves theoretically guaranteed unbiasedness but also exhibits superior empirical convergence, as shown in Fig. 1. Interestingly, we point out that not all explicit bias modeling schemes bring theoretically guaranteed unbiasedness, and we further generalize TranSUN into a novel generic regression model family, termed Generalized TranSUN (GTS), to shed light on the underlying mechanism to intrinsically maintain model’s unbiasedness. Essentially, GTS establishes the intrinsic unbiasedness through exploiting target transformation with conditional linearity based on particular point estimates [20] of $\mathrm{T(}y)|x$, exhibiting a more general model assumption compared to the vanilla MSE model. With this generic framework, not only various bias-free regression models can be directly devised to cope with varying data scenarios, but also any given regression model can be flexibly debiased in a plug-and-play manner even if the given model does not utilize any target transformation.
The major contributions can be summarized as follows.
  • We propose a novel preemptive paradigm to eradicate retransformation bias intrinsically from regression models to challenge the conventional ones, which is also the first work in the recommender system community to address the long-neglected bias problem.
  • We propose the novel TranSUN, a theoretically grounded, empirical convergence-friendly, and transformation-agnostic methodology, to endow the transformed MSE model with intrinsic unbiasedness exactly during the training phase.
  • We further propose a novel regression model family, termed Generalized TranSUN (GTS), to offer more theoretical insights, which also provides a generic framework for flexibly developing various bias-free regression models. Notably, our TranSUN and GTS have been validated to be directly applicable to other fields beyond the recommender system domain.
  • Extensive experiments on synthetic and real-world data validate the superiority and broad applicability of our methods. Now, TranSUN has been deployed in two core recommendation scenarios, i.e. product and short video recommendation, in Guess What You Like domain in the homepage of Taobao App (a leading e-commerce platform with DAU > 300M) to serve the major online traffic.
2 Related works
Regression models in recommender systems
mainly fall into two categories, i.e. value regression and classification-based regression, and our method belongs to the former. Value regression defines target as a continuous random variable and directly predicts target value using the expectation of the hypothetical distribution [22, 7, 48, 41, 42, 44], e.g. Gaussian distribution adopted by the MSE-based model family [42, 44, 43], ZILN distribution [40, 41]. By contrast, classification-based regression defines target as a discrete random variable via dividing it into multiple bins [45, 46, 9, 47], which mainly introduces either hierarchical routing [23, 21, 24, 39] or ordinal dependencies [23, 38] to boost models. Despite the progress, all the prior works neglect the study of model biasedness problems (e.g. retransformation bias problem), which probably lead to the systematic errors in the model predictions (as shown in Eq. (3)), e.g. systematic underestimation of duration models in video recommender systems. Differently, our paper specifically addresses this critical biasedness problem.
Retransformation bias
was initially identified in [2], with several works deriving theoretical correction terms by assuming the transformed distribution [13, 3, 8, 31]. Subsequent studies advanced correction frameworks free from distributional assumptions [12, 35, 5, 37, 27], achieving notable success across various fields [11, 19, 18, 30, 25, 1, 14, 34, 36, 33, 37]. Nonetheless, these approaches are still post-hoc remedies out of the model, bringing practical challenges when applied to industrial recommender systems. Differently, we preemptively address the bias via intrinsic model refinement.
Conditional transformation models (CTM)
learns a parametric transformation introducing interaction between target $y$ and input $x$ to transform them into variables of assumed distributions [16, 28, 4]. CLTM is further restricted to using affine transformation for enhanced interpretability [29], whose model assumption is $\alpha(x)+h_{0}(\mathrm{Y}_{x})\cdot\beta(x)\sim\mathcal{N}(0,1)$ with $\mathrm{Y}_{x}$ denoting $y|x$ and $h_{0}(\cdot)$ being parameter-free. Though CLTM and our GTS are superficially similar, they have one key differences, i.e. different dependencies in model assumptions, where the slope $\beta(x)$ in CLTM’s is only related to $x$, while ours (Eq. (11)) is related to both $x$ and $y$. In addition, regarding model classes and motivations, CLTM belongs to transformation models [17] for interval estimation, while GTS is a value regression model for unbiased point estimation.
3 Method
3.1 Preliminary
Given a training set $\mathcal{D}=\{(x,y)|(x,y)\sim\mathcal{P}(\mathrm{X,Y})\}$, where $x$ denotes the input features and $y$ denotes the continuous target label, the goal of the regression task is to make the model prediction $f(x;\theta)$ and the ground truth $y$ as close as possible, where $\theta$ is the parameter of the model $f$. In the recommender system community, a transformed MSE model is commonly adopted to boost the model convergence performance on high-skewed data, which is formulated as
$\mathcal{L}_{\mathrm{MSE}}^{\mathrm{T}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\big(f(x;\theta)-\mathrm{T}(y)\big)^{2}\Big],$ (2)
where $\mathrm{T}$ is a customized bijective mapping. However, since minimizing Eq. (2) makes $f(x;\theta)$ be an estimate of $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\big[\mathrm{T}(y)\big]$ (derived in Appendix B.1), directly back-transforming $f(x;\theta)$ leads to the retransformation bias problem, as shown in Eq. (1). Taking a common example in the industrial community, assuming the inverse transformation $\mathrm{T}^{-1}$ is convex, e.g. $\mathrm{T}(y)$ is $\log(y+1)$ or $\sqrt{y}$, then the retransformation bias will result in the systematic underestimation of model prediction $\hat{y}|x$, as
$\hat{y}|x=\mathrm{T}^{-1}(f(x;\theta))=\mathrm{T}^{-1}\Big(\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\big[\mathrm{T}(y)\big]\Big)\leq\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\Big[\mathrm{T}^{-1}\big(\mathrm{T}(y)\big)\Big]=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y],$ (3)
where $\leq$ is obtained by Jensen’s inequality and it is reversed when $\mathrm{T}^{-1}$ is concave.
3.2 TranSUN: joint bias learning
To intrinsically eradicate the retransformation bias, we propose TranSUN, which introduces an auxiliary branch $z(x;\theta_{z})$ to the transformed MSE model to explicitly learn the bias. Specifically, we employ a multiplicative bias modeling scheme, i.e. supervising $z(x;\theta_{z})$ to learn the ratio of the ground truth $y$ to the biased prediction $\mathrm{T}^{-1}\big(f(x;\theta)\big)$, hence the bias learning loss is formulated as
$\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\Big]\Big)^{2}\Big],$ (4)
where $\mathrm{stop\_grad}$ is applied to prevent the gradients of $\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}$ from passing to $\theta$, $|x|$ denotes the absolute value of $x$, and $\epsilon$ is a positive hyperparameter for avoiding division by zero. The reason for adopting this scheme is that the target ratio has intuitively small variance, empirically making the loss much smoother (as shown in Fig. 3a and discussed in Appendix A.2). During training, our model jointly learns the bias modeling task and the main regression task to achieve preemptive debiasing, thus the total loss is formulated as
$\mathcal{L}_{\mathrm{TranSUN}}^{\mathrm{T}}=\mathcal{L}_{\mathrm{MSE}}^{\mathrm{T}}+\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}.$ (5)
In the testing phase, the model prediction $\hat{y}|x$ is formulated as
$\hat{y}|x=z(x;\theta_{z})\cdot\big(|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon\big).$ (6)
Theoretical analysis of unbiasedness.
To illustrate the inherent unbiasedness of our method, we prove the prediction $\hat{y}|x$ to be an estimate of $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]$ as follows (see Appendix B.2 for details). Owing to the $\mathrm{stop\_grad}$ operation, the optimization of the total loss $\mathcal{L}_{\mathrm{TranSUN}}^{\mathrm{T}}$ can be decomposed into two independent parts, as
$\min_{\theta,\theta_{z}}\mathcal{L}_{\mathrm{TranSUN}}^{\mathrm{T}}\Leftrightarrow\min_{\theta}\mathcal{L}_{\mathrm{MSE}}^{\mathrm{T}}+\min_{\theta_{z}}\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}.$ (7)
In the second part, the prediction $\hat{y}|x$ is derived as the estimate of $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]$ via
$\min_{\theta_{z}}\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\,\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\Big[z(x;\theta_{z})-\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\Big]=0$ (8)
$\Leftrightarrow z(x;\theta_{z})\cdot\big(|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon\big)=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]\ \Leftrightarrow\ \hat{y}|x=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y].$
Key discussion.
According to Eq. (8), we propose an interesting argument: not all bias learning schemes provide theoretically guaranteed unbiasedness. As an example, supposing we turn the learning target of $z(x;\theta_{z})$ in Eq. (4) to the ratio of biased prediction to ground truth, i.e. ${\mathrm{T}^{-1}(f(x;\theta))}/{y}$, then the model prediction is derived as an estimate of ${1}/{\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[\frac{1}{y}]}$ (derived in Appendix B.3), which is a biased estimate bringing model underestimation according to Eq. (3). More empirical results are shown in Table 6. Therefore, in the next section, we point out that it is conditional linearity as the essential mechanism for the unbiasedness of TranSUN, rather than the explicit bias learning.
3.3 Generalized TranSUN (GTS)
In this section, we generalize TranSUN to a novel generic regression model family, termed Generalized TranSUN (GTS), to provide more insights into the intrinsic unbiasedness mechanism of our method. The total loss function of GTS is formulated as
$\mathcal{L}_{\mathrm{GTS}}=\mathcal{L}_{\mathcal{H}_{q}}\Big(x,\mathrm{T}(y);\theta_{q}\Big)+\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-y\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big]\Big)^{2}\Big],$ (9)
where $\mathcal{H}_{q}$ is a known apriori distribution parameterized by $x$ and $\theta_{q}$, $f(x;\theta_{q})$ is the expectation of the hypothesized distribution $\mathcal{H}_{q}(x,\theta_{q})$, and $\kappa$ is a parameter-free function with $\kappa(\cdot)>0$.
Specifically, the first loss term $\mathcal{L}_{\mathcal{H}_{q}}$ is termed conditional point loss, which supervises $f(x;\theta_{q})$ to specifically learn a conditional point estimate $\mathbb{Q}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\mathrm{T(}y)\big]$ (e.g. arithmetic mean, median, mode) via customizing the apriori conditional probability assumption of $\mathrm{T}(y)|x\sim\mathcal{H}_{q}(x,\theta_{q})$, with
$f(x;\theta_{q})=\mathbb{E}_{\mathrm{T}(y)\sim\mathcal{H}_{q}(x,\theta_{q})}[\mathrm{T}(y)]=\mathbb{Q}_{\mathrm{T}(y)\sim\mathcal{P}(\mathrm{T(Y)}|x)}\big[\mathrm{T(}y)\big]=\mathbb{Q}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\mathrm{T(}y)\big].$ (10)
The second loss term in Eq. (9) is termed linear transformation loss, which much resembles linear-transformed MSE, whereas the key difference is that our slope $\kappa\big(f(x;\theta_{q})\big)$ is dynamically generated by the learned conditional point $f(x;\theta_{q})$. Additionally, due to the $\mathrm{stop\_grad}$ operation, the slope $\kappa\big(f(x;\theta_{q})\big)$ is actually completely determined by $\mathcal{L}_{\mathcal{H}_{q}}$. In the testing phase, the prediction of GTS is $\hat{y}|x={z(x;\theta_{z})}/{\kappa\big(f(x;\theta_{q})\big)}$, which can also be proved unbiased by following the derivation process in Eq. (7, 8) (also derived in Appendix B.4).
Model assumption.
Given any sample $x$, GTS essentially exploits a customized $x$-conditioned point (abbr. $\mathbb{Q}$) to generate a $\mathbb{Q}$-conditioned slope $\kappa$ for linear transformation of target $y$. Hence, the model assumption of GTS is formulated as
$z(x;\theta_{z})+\mathrm{Y}_{x}\cdot\kappa\big(\mathbb{Q}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\mathrm{T(}y)\big]\big)\sim\mathcal{N}(0,\sigma^{2}),$ (11)
where $\mathrm{Y}_{x}=(y|x)$ and $\mathcal{N}$ denotes Gaussian distribution. As shown in Eq. (11), the slope $\kappa$ is conditioned on $\mathcal{P}({\mathrm{Y}|x)}$, making GTS akin to a piecewise-linear-transformed MSE where the intervals are determined by the condition $x$. Accordingly, GTS can be viewed as a conditional linear-transformed MSE, the unbiasedness of which is naturally maintained since linear-transformed MSE is unbiased itself. Notably the violation of model assumptions will not invalidate the model’s theoretical unbiasedness, which only affects the model performance (e.g. convergence). Furthermore, according to Eq. (11), the variance of the hypothetical distribution of $y|x$ is modeled as $\sigma^{2}/\kappa(\mathbb{Q})$, where one can explicitly model the heteroscedasticity by specifying $\mathbb{Q}$ and $\kappa$ with the domain knowledge, exhibiting a more general model assumption than the traditional MSE model.
Application.
There are two main application scenarios for the proposed GTS. The first is to directly devise various bias-free regression models by customizing the assumption $\mathrm{T}(y)|x\sim\mathcal{H}_{q}(x,\theta_{q})$ and $\kappa$’s function type to cope with different data scenarios. Some design guidelines are shown in Appendix A.6, with reference results provided in Table 1510. The second is to serve as a plug-and-play debiasing module to directly eradicate bias from any given regression model, even if the given model does not use any transformation $\mathrm{T}$. Specifically, for any given regression model $\mathcal{M}$, one can easily establish the debiased model $\mathcal{M}^{\prime}$ by setting $\mathrm{T}$ to identity transformation, $\mathcal{L}_{\mathcal{H}_{q}}$ to the loss function of $\mathcal{M}$, and $f(x;\theta_{q})$ to the prediction of $\mathcal{M}$. More empirical results are shown in Table 6.
4 Experiments on synthetic data
In this section, we conduct experiments on synthetic data to evaluate model performance. Specifically, we sample abundant instances $\{y_{i}\}_{i=1}^{N}$ from the pre-defined distribution $\mathcal{H}(\mathrm{Y})$ to form the conditional distribution $\mathcal{P}_{s}(\mathrm{Y}|x)$, and then models are trained on $\{y_{i}\}_{i=1}^{N}$ for learning regression task, after which the error between model prediction and the theoretical-calculated expectation $\mathbb{E}_{y\sim\mathcal{H}\mathrm{(Y)}}[y]$ are computed for evaluation.
4.1 Experiment settings
Figure 2: Density and statistics of the sampled $\mathcal{P}_{s}(\mathrm{Y}|x)$ from the eight pre-defined $\mathcal{H}(\mathrm{Y})$.
Data and metric.
We adopt $8$ distributions $\mathcal{H}(\mathrm{Y})$ from three categories: right-skewed, left-skewed, and symmetric. Details are shown in Appendix D.1.1. Fig. 2 illustrates the density and some statistics of the sampled $\mathcal{P}_{s}(\mathrm{Y}|x)$. We adopt SRE (Signed Relative Error) as the metric, i.e. ${(\hat{y}-y)}/{y}$ where $\hat{y}$ and $y$ are prediction and ground truth, respectively.
Implementation details.
See Appendix D.1.2 for the detailed introduction.
4.2 Results on synthetic data
Effectiveness on eradicating retransformation bias.
Three transformations, i.e. linear, square, and logarithmic transformations, are employed to compare the performance between transformed MSE and our TranSUN. As demonstrated in Table 1, the retransformation bias problem arises in transformed MSE when the transformation is not linear, i.e. model becoming underestimated with $\mathrm{T}(y)=\ln(y+1)$ and overestimated with $\mathrm{T}(y)=y^{2}$, which is consistent with Eq. (3). In contrast, TranSUN maintains unbiasedness (||SRE|| < 0.7%) regardless of the transformation applied. Notably MSE and linear-transformed MSE both exhibit unbiasedness in Table 1, that is because they converge well due to the extreme simplicity of input features and model architectures. Differently, MSE exhibits significantly poor convergence on real-world data, as detailed in the next section (Table 78).
Table 1: Comparison results of transformed MSE (T-MSE) and TranSUN under different transformations $\mathrm{T}$ on synthetic data. Red boldface indicates that |SRE| >1%. The difference to T-MSE is statistically significant at 0.02 level.
Methods $\mathrm{T}(\cdot)$ RS-G RS-BU RS-ZIG LS-B LS-BU SM-U SM-TN SM-BU
MSE - 0.0010 0.0023 0.0015 -0.0013 -0.0022 -0.0065 0.0005 -0.0043
T-MSE -0.0012 0.0023 -0.0013 -0.0017 -0.0024 -0.0065 -0.0003 -0.0042
TranSUN $\mathrm{T}(y)=0.5y$ -0.0011 0.0022 -0.0014 -0.0015 -0.0020 -0.0066 -0.0006 -0.0045
T-MSE -0.1401 -0.5108 -0.4481 -0.0179 -0.1587 -0.2447 -0.0217 -0.5341
TranSUN $\mathrm{T}(y)=\ln(y+1)$ 0.0014 0.0013 -0.0010 0.0014 0.0033 -0.0038 -0.0017 -0.0058
T-MSE 0.2122 1.1015 1.7534 0.0479 0.0506 0.1442 0.0193 0.3264
TranSUN $\mathrm{T}(y)=y^{2}$ -0.0045 0.0038 0.0019 0.0016 0.0063 -0.0069 -0.0035 0.0025
[…]
Methods RS-G RS-BU RS-ZIG LS-B LS-BU SM-U SM-TN SM-BU
[…]
WLR [9] - - 0.2483 - - - - -
[…]
MDME [21] -0.1132 -0.5876 -0.1758 0.2730 0.1004 0.5500 -0.1146 -0.7511
[…]
Table 3: Comparison of transformed MSE (T-MSE) and TranSUN under different transformations $\mathrm{T}$ on CIKM16 train and DTMart train. Bold indicates that gain <-70%. The differences are statistically significant at 0.05 level.
Methods $\mathrm{T}(\cdot)$ CIKM16 train DTMart train
TRE $\downarrow$
MRE $\downarrow$
TRE $\downarrow$
MRE $\downarrow$
T-MSE $\mathrm{T}(y)=\ln(y+1)$ 0.3468 0.3352 0.2894 0.1432
TranSUN 0.0133 (-96.2%) 0.0171 (-94.9%) 0.0803 (-72.3%) 0.0725 (-49.4%)
T-MSE $\mathrm{T}(y)=\sqrt{y}$ 0.1386 0.1243 0.4388 0.3421
TranSUN 0.0388 (-72.0%) 0.0283 (-77.2%) 0.0907 (-79.3%) 0.066 (-80.7%)
T-MSE $\mathrm{T}(y)=\arctan(y)$ 3.6266 10.9861 0.6936 0.3678
TranSUN 0.3528 (-90.3%) 1.2674 (-88.5%) 0.1554 ( -77.6%) 0.0678 (-81.6%)
Comparison with state-of-the-arts.
Our TranSUN is further compared with state-of-the-art regression models in the recommender system field. As shown in Table 2, our TranSUN achieves the best performance as it is the only method that maintains unbiasedness across all distributions. Subsequently, we provide an analysis of the biasedness of the baselines. The MAE model overestimates the mean on left-skewed and underestimates it on right-skewed data due to its median modeling approach. WLR exhibits an error of $0.2483$ on RS-ZIG, roughly equivalent to ${p_{+}}/{(1-p_{+})}=0.25$ ($p_{+}$} is the positive rate of RS-ZIG), consistent with the derivations in [9]. ZILN and OptDist both show significant biasedness on certain distributions, owing to these severely violating the LogNormal or ZILN distribution assumptions. MDME overestimates the mean on left-skew and underestimates it on right-skew data, likely because it estimates the mean of mode buckets under mode sub-distribution, which also explains its biasedness on SM-BU and SM-U as sampling error may result in severe mode estimation error on these data. TPM underestimates the mean on left-skew and overestimates it on right-skew data, likely because it learns the average of quantiles. Note that increasing bin number can reduce its bias, as shown in Sec E.1. CREAD presents a similar bias problem to TPM, whereas slightly alleviated due to its EAD binning method [38].
Comparison of GTS instance models.
Some common $\mathcal{L}_{\mathcal{H}_{q}}$ and $\kappa$ are selected to devise different instances of GTS, providing reference results for its application. As exhibited in Table 15, in terms of $\mathcal{L}_{\mathcal{H}_{q}}$, MSE loss is empirically the most versatile loss as it fastly converges to the unbiased solution in all distributions, whose $\mathbb{Q}$ corresponds to the arithmetic mean point. Among other losses, MAE loss presents better convergence on right-skewed data, while MSPE and MAPE loss both converge more effectively on left-skewed and symmetric data. Regarding function $\kappa$, all candidates exhibit unbiasedness consistent with TranSUN. The performance of $\kappa$ on real-world data is shown in Table 10.
5 Experiments on real-world data
5.1 Experiment settings
Datasets and metrics.
[…]
Implementation details.
See Appendix D.2.3 for the detailed introduction. TensorFlow-style implementation of our TranSUN and GTS is exhibited in Code 1. Some optional techniques for optimizing TranSUN’s online performance are introduced in Appendix A.4. Moreover, the efficiency analysis of our method is demonstrated in Appendix A.3.
5.2 Results
Effectiveness on eradicating retransformation bias.
Since these target data are right-skewed, commonly-used concave functions, i.e. logarithm, square, and arctan transformation, are adopted for performance comparison between transformed MSE and our TranSUN. Because retransformation bias can be viewed as fitting deviation, evaluations are conducted on the training sets of CIKM16 and DTMart. As shown in Table 3, transformed MSE exhibits significant bias issues, with $\arctan$ transformation being the most severe. Differently, TranSUN significantly mitigates the bias (mostly < -70.0%) for all transformations. Furthermore, we assessed their performance on different target bins on Indus, where bins are obtained by dividing ascending targets into equal frequencies. In Fig. 3c, LogMSE and LogSUN denotes transformed MSE and TranSUN with $\mathrm{T}(y)=\ln(y+1)$, respectively, with Signed TRE used as metric. As illustrated, LogSUN substantially alleviated bias across all bins compared to LogMSE. Moreover, LogMSE’s bias issue gets more severe when the target value grows larger, and meanwhile, our LogSUN’s debiasing intensity gets higher as well. Fig. 3c further shows $\kappa$ gradually decreases as target $y$ increases, to provide relatively stable $\kappa\cdot y$ for smoothing $\mathcal{L}_{\mathrm{sun}}^{\ln}$.
Comparison with post-hoc correction methods.
To further demonstrate the debiasing capability of our method, our LogSUN is compared against two classic bias correction methods, i.e. Normal Theory Estimates (NTE) [11] and Smearing [12], alongside a mainstream calibration method, i.e. SIR [10], on CIKM16 test. Different from LogSUN, all the baselines are post-hoc methods, which naturally confer them an advantage in the comparison. Nevertheless, as shown in Table 6, our method achieves comparable performance with these baselines. Among them, the poor performance of NTE is attributed to the transformed target not adhering to the Gaussian distribution assumptions. In contrast, Smearing and SIR exhibit superior results, as they are free from distributional assumptions.
[…]
Models Correction methods TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$
LogMSE - 0.3451 0.3667 0.6189 0.4528
LogMSE NTE [11] 0.1262 0.1442 0.5718 0.4369
[…]
LogMSE SIR [10] 0.0198 0.0236 0.5821 0.4309
[…]
Table 5: Comparison of LogSUN’s different bias modeling schemes on CIKM16 test, where bold indicates the best.
Scheme ID $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}}$ TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
S0 $\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[y-\exp(f(x;\theta))\Big]\Big)^{2}$ 0.0689 0.0829 0.6153 0.4459 0.6711
S1 $\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[{\exp(f(x;\theta))}/{y}\Big]\Big)^{2}$ 0.1288 0.1036 0.6255 0.4387 0.6730
S2 (Ours) $\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[y\cdot\exp(-f(x;\theta))\Big]\Big)^{2}$ 0.0123 0.0439 0.5625 0.4333 0.6740
Table 6: Results of extending our method to advanced regression models on CIKM16 test. Except for XAUC, metric gains are statistically significant at 0.05 level.
Extensions TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
WLR [9] + GTS 0.0286 (-88.72%) 0.0714 (-78.36%) 0.5814 (-33.20%) 0.4735 (-22.42%) 0.6737 (-0.07%)
ZILN [40] + GTS 0.0250 (-94.31%) 0.0223 (-95.04%) 0.5580(-12.81%) 0.4308 (-7.53%) 0.6745 (0.01%)
MDME [21] + GTS 0.0828 (-81.38%) 0.0254 (-73.98%) 1.0458 (-33.60%) 0.7874 (-30.85%) 0.6743 (0.19%)
OptDist [41] + GTS 0.0234 (-93.88%) 0.0248 (-94.96%) 0.5604 (-10.84%) 0.4375 (-4.66%) 0.6748 (-0.09%)
[…]
Models CIKM16 test DTMart test
TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$ TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
MSE 0.1139 0.1687 0.6192 0.5024 0.6693 0.1028 0.1349 0.4893 0.2579 0.9369
MAE 0.0819 0.1281 0.6837 0.4285 0.6747 0.1655 0.0917 0.4940 0.2237 0.9388
LogMSE 0.3451 0.3667 0.6189 0.4528 0.6745 0.2889 0.1448 0.5232 0.2494 0.9440
[…]
WLR [9] 0.2536 0.3299 0.8703 0.6103 0.6742 0.0916 0.7488 1.0178 0.5770 0.9249
[…]
MDME [21] 0.4448 0.0976 1.5749 1.1387 0.6730 0.2149 0.0861 0.9654 0.3395 0.9296
[…]
Table 8: Comparison of advanced regression methods on Indus test, with bold indicating the best.
Indus test Indus Top test
Models TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$ TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
MSE 0.3253 1.6568 2.4540 0.6018 0.7737 0.7719 2.5136 1.6605 0.5845 0.6703
LogMSE 0.6080 0.7463 2.4105 0.5871 0.7819 1.1334 2.4296 1.6130 0.5475 0.6593
LogMAE 0.4685 0.8314 2.4394 0.5691 0.7816 0.8973 2.5427 1.6427 0.5273 0.6567
WLR [9] 0.6940 0.7399 2.5622 0.6419 0.7751 0.7883 1.3097 1.6403 0.5673 0.6608
ZILN [40] 0.2945 0.4159 2.6586 0.5890 0.7809 0.7547 1.5717 1.5543 0.5277 0.6661
MDME [21] 0.3987 0.4585 2.4880 0.6225 0.7790 0.7986 1.5056 1.5789 0.5494 0.6677
TPM [23] 0.0652 0.1350 2.3981 0.6186 0.7799 0.4218 0.9261 1.5706 0.5052 0.6730
CREAD [38] 0.0917 0.1180 2.3353 0.5815 0.7808 0.6484 1.0733 1.5825 0.5191 0.6693
OptDist [41] 0.2867 0.3650 2.3486 0.5780 0.7822 0.7247 1.4571 1.5614 0.5212 0.6679
LogSUN 0.0197 0.0410 2.2765 0.5683 0.7815 0.3730 0.8602 1.4969 0.4801 0.6749
Performance comparison with state-of-the-arts.
Our TranSUN is compared against common regression loss (e.g. MSE, MAE), transformed models (e.g. LogMSE, SqrtMSE), advanced classification-based regression models (e.g. TPM, MDME), and fancy value regression models (e.g. ZILN, OptDist) on CIKM16 test and DTMart test. In Table 7, LogSUN and SqrtSUN denote our method with $\mathrm{T}(y)$ set to $\ln(y+1)$ and $\sqrt{y}$, respectively. As illustrated, our method achieves the best performance across our TRE and MRE metrics, as well as the commonly used NRMSE and NMAE metrics. Note that applying TranSUN usually results in a slight decline in XAUC. However, this trade-off remains favorable as discussed in Appendix A.1. Furthermore, as shown in Table 8, consistent superiority of our LogSUN is observed on Indus test, where its XAUC for top high-priced samples also achieves the best performance (> +0.015 uplift to LogMSE).
Extension to state-of-the-arts.
To validate the effectiveness of GTS as a debiasing framework, we apply our method to several state-of-the-art models with significant bias issues on CIKM16 test, where $\kappa(f(x;\theta_{q}))=1/(f(x;\theta_{q})+1)$ is adopted in Eq. (9). As depicted in Table 6, the bias issues of classification-based models (i.e. MDME, WLR) and value regression models (i.e. ZILN, OptDist) are all significantly mitigated (< -70.0% in TRE and MRE) by integrating our approach.
Table 9: Results of the online A/B test, with the 95% confidence interval (CI-95%) and $p$-value.
Scenarios Product recommendation Short video recommendation
A/B tests LogSUN (Ours) vs. LogMAE (Base) LogSUN (Ours) vs. LogMSE LogSUN (Ours) vs. LogMAE (Base)
Traffics & periods 15% vs. 20% for 1 week 15% vs. 9% for 1 week 6% vs. 6% for 1 month
Metrics GMV GMVPI GMV GMVPI GMV GMVPI
Gains +1.06% +0.92% +1.02% +0.75% +1.83% +1.57%
95%CI [ 0.52% , 1.59% ] [ 0.32% , 1.51% ] [ 0.51% , 1.54% ] [ 0.18% , 1.32% ] [ 0.50% , 3.17% ] [ 0.24% , 2.89% ]
pp-value 0.0001 0.0024 0.0001 0.0096 0.0065 0.0198
Online A/B test.
In Taobao App (DAU > 300M), our LogSUN is deployed to serve as a pointwise ranking model for GMV target in two recommendation scenarios (i.e. product and short video recommendation in Guess What You Like business domain in Taobao App homepage), respectively, to optimize object distribution (see Appendix E.2 for details), which is compared against the previous online base (LogMAE) and LogMSE. Two common metrics, GMV and GMVPI (GMV Per Impression), are adopted to assess online performance. Due to the large variability of these GMV-related metrics, we allocated either substantial traffic (e.g. 15% of the total traffic for LogSUN) or a sufficiently long period (e.g. one month) for A/B tests to obtain statistically significant results. As shown in Table 9, LogSUN achieves significant gains (e.g. > +1.0% GMV gains) over all the baselines. Now, LogSUN has been deployed to serve the major traffic.
5.3 Ablation studies
(a) Loss curves
(b) Metric curves
(c) Binned STRE & $\kappa$
(d) Impact of $\epsilon$
Figure 3: Visualization results across experiments. (a) Training loss curves for LogSUN with different schemes on CIKM16. (b) Training curves of TRE and MRE for LogSUN with different schemes on CIKM16. (c) Signed TRE (STRE) and average slope $\bar{\kappa}$ on divided target bins of Indus test. (d) The sensitivity of LogSUN to $\epsilon$ on CIKM16 test.
Comparison of different GTS instance models.
To offer reference results for GTS’s application on real-world data, some common $\mathcal{L}_{\mathcal{H}_{q}}$ and $\kappa$ are selected to build different instances of GTS for comparison. As shown in Table 10, MSE loss is the best-performing $\mathcal{L}_{\mathcal{H}_{q}}$, while MAE loss also performs well as the target data are right-skewed, which is consistent with the results in Table 15. Regarding function $\kappa$, compared to the one used in TranSUN, other $\kappa$ candidates all bring about poorer loss convergence, which is probably because the variances of their $\kappa\cdot y$ are all too large (shown in Table 17), leading to an excessively steep optimization landscape for $\theta_{z}$.
Comparison of different bias modeling schemes.
To show the superiority of the multiplicative bias modeling scheme in Eq. (4), two other intuitive schemes are compared with ours on the LogSUN model, as presented in Table 6, where notably neither of the scheme candidates (S0, S1) belongs to GTS family. The impact of different schemes on model training is illustrated in Fig. 3a, 3b. As visible, despite theoretically unbiased, S0 scheme exhibits poor convergence performance, likely due to the high loss variance caused by its additive modeling scheme. S1 scheme converges well but suffers from significant bias, as it is not theoretically unbiased, as discussed in Sec. 3.2. In contrast, our scheme (S2) achieves both superior convergence and theoretical unbiasedness.
[…]
$\mathcal{L}_{\mathcal{H}_{q}}$ $\kappa(\cdot)$ TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
[…]
MSPE $\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.1077 0.0679 0.5856 0.4595 0.6731
[…]
MSE $\kappa(u)=1/(|u|+\epsilon)$ 0.0399 0.0680 0.5898 0.4753 0.6708
MSE $\kappa(u)=|u|$ 0.1810 0.2483 0.6328 0.5423 0.6633
Architectural sensitivity.
Since our method introduces an extra bias learning branch to transformed MSE, we explore the impact of different parameter sharing schemes between the added branch and the original model, which is conducted on the LogSUN model. In Table 16, sharing parameters from bottom layers (e.g. embedding layers) has minimal impact on performance, whereas further sharing more parameters within MLP results in a noticeable performance degradation, which might be inferred that the added task primarily affects the downstream MLP-modeled fitting function rather than the distribution of latent embedding.
Impact of hyperparameters.
We study the impact of the only hyperparameter $\epsilon$ on the LogSUN model using CIKM16. As visible in Fig. 3d, LogSUN stably demonstrates superior performance when $\epsilon$ is between $0.01$ and $1.0$. As $\epsilon$ further approaches zero, the performance rapidly degrades. This may be because $\mathrm{T}^{-1}(f(x;\theta))$ tends to be near zero at the early stages of training, thus ${y}/{(|\mathrm{T}^{-1}(f(x;\theta))|+\epsilon)}$ gets extremely large when $\epsilon$ is close to zero as well, causing significant loss oscillation and hindering model convergence. Conversely, performance also degrades when $\epsilon\geq 100$, likely due to ${y}/{(|\mathrm{T}^{-1}(f(x;\theta))|+\epsilon)}$ being too small to be learned within the fitting scope of $z(x;\theta_{z})$, thus leading to the model underfitting.
6 Conclusion
In this paper, we propose a novel preemptive paradigm to eradicate retransformation bias intrinsically from the regression models to challenge the conventional ones, as the first work in the recommender system field to address the long-neglected bias problem. Specifically, a novel method, TranSUN, is proposed with a joint bias learning manner to offer theoretically guaranteed unbiasedness under empirical superior convergence. It is further generalized into a novel generic regression model family, Generalized TranSUN (GTS), which not only offers theoretical insights into TranSUN but also serves as a generic framework for flexibly developing various bias-free models. Extensive experiments demonstrate the superiority and broad applicability of our methods.
Limitations and future works.
Regarding real-world applications, although the effectiveness of TranSUN has been validated across data from various domains, these target distributions are all right-skewed. Hence, we expect more exploration for its application where the target distributions are not right-skewed. Moreover, a limitation of our method is the decline in XAUC, which conforms to the No Free Lunch (as discussed in Appendix A.1), thus integrating isotonicity into TranSUN is a challenging but interesting direction. In the future, we plan to investigate GTS further, e.g. exploring a general methodology to identify the optimal point $\mathbb{Q}$ and function $\kappa$ on any given data.
Acknowledgments
We deeply appreciate Han Wu, Chi Li, Kun Yuan, Runfeng Zhang, Weiqiu Wang, Longbin Li, Xin Yao, Nan Hu, Chuan Wang, and Hanqi Jin for their insightful suggestions and discussions. We thank Kewei Zhu, Yuxin Duan, Hao Fang, and Gaofeng Li for necessary supports about online serving. We thank anonymous reviewers for their constructive feedback and helpful comments.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes. Our main contributions are also detailed in Sec. 1. Also see Sec. 3, Appendix B for more theoretical evidence and Sec. 45 for more experimental evidence.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes. We discussed the limitations of our work in Sec. 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We detail the model assumption in Sec. 3.3 and proof of theoretical unbiasedness in Sec. 3.2, Appendix B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Regarding the data, the codes for producing the synthetic data is presented in Code 2, and for real-world ones two publicly-accessable datasets (CIKM16, DTMart) are adopted for reproducibility, as detailed in Appendix D.1.1D.2.1. Regarding the implementation, all experimental details for synthetic and real-world data are presented in Appendix D.1.2D.2.3, respectively.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Regarding the data, we provide the codes for producing the synthetic data in Code 2, and for the real-world data scenario, we adopt two publicly-accessible datasets (CIKM16, DTMart). And we’ll open-source our code of our method implementation.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We detail the experimental settings for synthetic and real-world data in Appendix D.1.2D.2.3, respectively.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report statistical significance in the key results, including Table 1396.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include the related information in Appendix D.1.2D.2.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the NeurIPS Code of Ethics and made sure that the paper conforms to it.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: As discussed in Appendix F, though our method has been applied in two core recommendation scenarios on a leading e-commerce platform (DAU > 300M) to serve the major online traffic, there are no obvious paths that lead to potential harm since our work just present a unbiased regression methodology and is not tied to particular applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not foresee any high risk for misuse of this work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, we credited them in appropriate ways.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Frequently Asked Questions (FAQ)
A.1 Root reason for the trade-off between point accuracy (e.g. NRMSE) and XAUC
Here we elaborate on the root cause of the trade-off phenomenon between point accuracy and XAUC, and further highlight the limitations of XAUC. This limitation leads to the inconsistency between the offline XAUC metric and online business metrics (as shown in Table 9), making XAUC no longer suitable for offline assessment. To address this, we supplemented the offline evaluation with two more consistent offline metrics, high-value XAUC (as shown in Table 8) and a commonly used ranking metric NDCG (as shown in Table 11), on both of which our model demonstrates superior performance. For clarity, let us start with a concrete case as follows.
An illustrative case for trade-off:
We first sample two inputs $\{x_{1},x_{2}\}\sim\mathcal{P}(\mathrm{X})$, and then sample $3$ groups of target $y$ from $\mathcal{P}(\mathrm{Y}|x_{1})$ and $\mathcal{P}(\mathrm{Y}|x_{2})$, respectively, and let the samples be $\{3,3,3\}\sim\mathcal{P}(\mathrm{Y}|x_{1})$ and $\{8,1,1\}\sim\mathcal{P}(\mathrm{Y}|x_{2})$, and let $\mathrm{T}(x)=\log(x)$. Then, 1) The LogMSE prediction is $f(x_{1})=\mathrm{T}^{-1}(\mathbb{E}[\mathrm{T}(\mathrm{Y})|x_{1}])=\exp((\log(3)+\log(3)+\log(3))/3)=3$, $f(x_{2})=\mathrm{T}^{-1}(\mathbb{E}[\mathrm{T}(\mathrm{Y})|x_{2}])=\exp((\log(8)+\log(1)+\log(1))/3)=2$, where $f(x_{1})>f(x_{2})$ and the bias occurs on the sample $\mathrm{X}=x_{2}$; 2) The TranSUN/GTS prediction is $f(x_{1})=\mathbb{E}[\mathrm{Y}|x_{1}]=(3+3+3)/3=3,\,f(x_{2})=\mathbb{E}[\mathrm{Y}|x_{2}]=(1+1+8)/3=10/3$, where $f(x_{1})<f(x_{2})$; 3) The descending ranking result of LogMSE is $\{3,3,3,8,1,1\}$, with XAUC = 0.333, NRMSE = 0.324, and TRE = 4/15, while the result of TranSUN/GTS is $\{8,1,1,3,3,3\}$, with XAUC = 0.233, NRMSE = 0.301, and TRE = 0; 4) It is observed that a trade-off occurs between point accuracy (NRMSE, TRE) and XAUC, i.e. TranSUN/GTS achieves better point accuracy (NRMSE, TRE) while XAUC worsens.
Root cause of the trade-off:
1) Reason for the improvement of high-value XAUC & overall point accuracy metric: TranSUN/GTS compensates underestimation brought by retransformation bias, resulting in better point and rank accuracy of high-value samples (e.g. for sample $(x_{2},8)$, relative error improves from 0.75 to 0.58, and its rank order moves from 4th to 1st, consistent with results on Indus Top (Table 8)), which significantly improves the overall point accuracy (e.g. improvements in NRMSE & TRE in the above case) since these metrics are value-aware (i.e. metric calculation aware of absolute value of sample label); 2) Reason for the drop of overall XAUC: Correcting retransformation bias also leads to some hard low-value samples being overestimated (e.g. predictions for the two samples $(x_{2},1)$ increase from 2 to 10/3, consistent with Fig. 3c), which dominates the overall XAUC drop since there are far more hard low-value samples than high-value ones in highly right-skewed data and XAUC metric is value-agnostic (i.e. metric calculation only aware of the relative rank order of sample label but not aware to its absolute value).
Overall XAUC is inconsistent with online metrics, thus we use high-value XAUC and NDCG:
1) The core goal of a recommender system is to improve online business metrics, while XAUC is merely an intermediate metric. For business target modeled by regression (GMV, watch time, etc.), the event of ranking high-valued samples in the correct position is far more important for online metrics than that of low-valued samples. However, these two events are equally weighted (i.e. value-agnostic) in the calculation of XAUC, thereby being dominated by abundant low-valued samples when data are highly right-skewed, which probably contribute only a small portion of online metrics (e.g. 70% of the top low-valued samples contribute only 20% of GMV in Indus). Therefore, overall XAUC is no longer consistent with online metrics; 2) To address this, we supplemented the offline assessment with a more consistent metric, high-value XAUC (Table 8), based on which we achieved improved online results (Table 9). Additionally, NDCG also addresses XAUC’s limitation, so we also provide their results. As shown in Table 11, NDCG is computed within all samples (NDCG@All) and the top 10% high-value ones (NDCG@10%), respectively. Key observations include: a) TranSUN performs better in NDCG@All than in XAUC since NDCG is more robust to low-value samples. b) TranSUN shows superior performance in NDCG@10%, consistent with the reasons described in the 2nd point.
Table 11: NDCG metric results of our method and advanced baselines on CIKM16 test and DTMart test. The best results are in boldface.
Model CIKM16 test DTMart test
NDCG@All NDCG@10% NDCG@All NDCG@10%
MSE 0.9180 0.5201 0.9770 0.9692
MAE 0.9497 0.5654 0.9867 0.9665
LogMSE 0.9498 0.5517 0.9960 0.9736
WLR [9] 0.9475 0.5433 0.9693 0.9609
ZILN [40] 0.9485 0.5781 0.9963 0.9756
MDME [21] 0.9391 0.5671 0.9763 0.9639
TPM [23] 0.9502 0.5888 0.9960 0.9772
CREAD [38] 0.9504 0.5881 0.9962 0.9774
OptDist [41] 0.9500 0.5852 0.9965 0.9770
LogSUN 0.9499 0.5904 0.9967 0.9779
A.2 Convergence analysis for $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}},\,\mathcal{L}^{\mathrm{T}}_{\mathrm{MSE}},$ and $\mathcal{L}_{\mathrm{MSE}}$
As for convergence rate bound, since their loss formulations are all squared errors, the theoretical convergence rate bound using SGD is $O(\frac{1}{\sqrt{K}})$ under the Lipschitz assumption, where $K$ is update iteration number, according to [6]. The key difference in the convergence performance of $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}},\,\mathcal{L}^{\mathrm{T}}_{\mathrm{MSE}},$ and $\mathcal{L}_{\mathrm{MSE}}$ lies in their smoothness. This is because the variance and the outlier proportion of the target are different, resulting in optimization landscapes with different degrees of smoothness. As shown in Table 12, we present the target variance (Var), the variance divided by squared expectation (V/E2), and the outlier proportion (Outlier%) outside of 3 times the standard deviation of the three losses, where $\mathrm{T}(x)=\log(x)$. As shown, $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}}$ has a smaller variance and a better outlier distribution than the vanilla MSE, thus $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}}$ converges more smoothly (Fig.1). Therefore, TranSUN/GTS essentially decomposes a difficult task into two easier and smoother tasks.
Table 12: Convergence analysis for $\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}},\,\mathcal{L}^{\mathrm{T}}_{\mathrm{MSE}},$ and $\mathcal{L}_{\mathrm{MSE}}$
Losses CIKM16 DTMart
Var V/E2 Outlier% Var V/E2 Outlier%
$\mathcal{L}_{\mathrm{MSE}}$ 16440 0.41 1.03% 97494 12.1 0.78%
$\mathcal{L}^{\mathrm{\ln}}_{\mathrm{sun}}$ 0.63 0.36 0.59% 2.327 0.93 0.32%
$\mathcal{L}^{\ln}_{\mathrm{MSE}}$ 0.62 0.02 0.29% 1.87 0.15 0.28%
A.3 Efficiency analysis for the proposed TranSUN
As different parameter sharing architectures of TranSUN are explored in Appendix E.4, here we supplement the results (Table 16) with TranSUN’s additional memory (Mem) and computational overhead (MFLOPS), and calculate the relative increase (OH) compared with T-MSE. As shown in Table 13, the only-sharing-embedding architecture scheme is the most cost-effective, maintaining optimal performance while introducing much little overhead. On the other hand, we provide additional information on the resource overhead associated with the online TranSUN model in the e-commerce platform’s recommendation system, which achieves a high ROI measure (with Table 9): 1) Training: Under the same resources (800 workers), the update iteration (global step) number per second of LogMSE and TranSUN were 89.2 and 86.8, respectively. Therefore, applying TranSUN only reduces the training speed by 2.7%, with the training time increased by 2.78%; 2) Inference: The mean and p99 latency of the LogMSE were 27.8ms and 49.8ms, respectively, while the mean and p99 latency of TranSUN were 28.1ms and 51.6ms, respectively; 3) Deployment: In the online service with 300 million daily active users, deploying TranSUN did not increase CPU resources and only increased GPU resources by 5%.
Table 13: Comparison results of the LogSUN model under different parameter sharing architectures on CIKM16 test. “X/YX/Y” denotes sharing the first $X$-th layers of the network with totally $Y$ layers. The results with the best ROI are in boldface.
Parameter sharing schemes Mem (MB) OH (%) MFLOPS OH (%)
0/1 Embedding + 0/3 MLP 610.46 100.00 350.86 100.00
1/1 Embedding + 0/3 MLP 305.29 0.02 190.52 8.60
1/1 Embedding + 1/3 MLP 305.27 0.01 186.12 6.09
1/1 Embedding + 2/3 MLP 305.24 0.00 177.62 1.25
1/1 Embedding + 3/3 MLP 305.23 0.00 175.47 0.02
A.4 Optional techniques for optimizing LogSUN’s online performance
As mentioned in Appendix A.1, TranSUN likely results in a certain decrease in the overall XAUC, particularly for low-priced orders. Although these orders only contribute 20% of the total GMV, they represent the majority in view of the total number of transaction orders. Therefore, the reduced ranking accuracy of these orders’ GMV estimates deteriorates the accuracy of their overall ranking score, which in turn deteriorates the online metrics regarding the number of transaction orders. Therefore, it is crucial to lift the ranking accuracy of these samples under TranSUN’s modeling. Here are two techniques that have been validated as effective in our scenarios: 1) Selecting a suitable combination of ranking losses: As shown in lines 77-79 of the Code 1, selecting these two ranking losses carefully has been shown to improve the overall XAUC by > +0.01 on our private industrial data. 2) Fusion with the original predictions: As shown in lines 51-52 of the Code 1, a pre-defined fusion threshold can be used to restrict TranSUN’s functioning scope to high-priced samples, which improved the overall XAUC by > +0.003 on our private data. To select the threshold, we first divided the sample into 50 bins with equal frequency based on their original prediction values in ascending order. Then, we calculated the XAUC results of the original and de-biased prediction values within each bin, respectively. We found that as the bin number increased, the XAUC gain of de-biasing gradually shifted from negative to positive. Therefore, the left boundary of the bin that first achieved a positive XAUC gain was adopted as the fusion threshold. Once this threshold is determined, it does not need to be adjusted anymore, unless the data distribution shifts significantly.
A.5 Comparing GTS with probabilistic modeling methods (PMM), e.g. Mixture Density Networks (MDN), Normalizing Flows (NF)
When our GTS serves as a point estimator, it has three advantages over PMM: 1) Flexibility: It can be flexibly plugged into any model with minor revision (as validated in Table 6). Another advantage is that in industrial practice TranSUN can be fully warmed-up from the online base model, further accelerating TranSUN’s convergence; 2) Lightweight: It requires much less resource than PMM. For example, MDN’s overhead grows linearly with mixture components, and NF incurs significant overhead due to Jacobian computations; 3) Training stability: PMM is much harder to converge than GTS given their complex assumptions. For example, on Indus, OptDist [41] (an MDN case) needs 2.13x the iterations TranSUN requires to reach batch-NRMSE=$3.0$.
In addition, with minor revision to GTS’s model assumption (i.e. modeling constant $\sigma$ in Eq. (11) as a function of $x$), GTS can also directly model the conditional probability $\mathcal{P}(\mathrm{Y}|x)$ in the CTM manner [16]. Specifically, GTS’s assumption formulation turns to
$z(x;\theta_{z})+\mathrm{Y}_{x}\cdot s(x;\theta_{s})\cdot\kappa\big(\mathbb{Q}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\mathrm{T(}y)\big]\big)\sim\mathcal{N}(0,1),$ (12)
where $s(x;\theta_{s})>0$ actually models the reciprocal of the standard deviation $\sigma(x)$. In training stage, according to [29, 16], one can establish the scoring-rule-based loss $\mathcal{L}_{sr}(x,y;\mathbf{h})$ as
$\mathcal{L}_{sr}(x,y;\mathbf{h})=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\sum_{i=1}^{N}{\mathbf{1}\{y\leq v_{i}\}\log(\Phi(\mathbf{h}(x,v_{i})))+\mathbf{1}\{y>v_{i}\}\log(1-\Phi(\mathbf{h}(x,v_{i})))}\Big],$ (13)
where the $\mathbf{1}\{\cdot\}$ is the indicator function, $\{v_{i}\}_{i=1}^{N}$ denotes pre-defined ascending grid points of $y$, $\Phi(u)=\frac{1}{2}[1+\mathrm{erf}(\frac{u}{\sqrt{2}})]$, and $\mathbf{h}(x,y)$ is defined as
$\mathbf{h}(x,y)=-z(x;\theta_{z})+y\cdot s(x;\theta_{s})\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big].$ (14)
Therefore, the total loss is formulated as
$\mathcal{L}_{\mathrm{GTS}}^{\mathrm{pmm}}=\mathcal{L}_{\mathcal{H}_{q}}\big(x,\mathrm{T}(y);\theta_{q}\big)+\mathcal{L}_{sr}(x,y;\mathbf{h}).$ (15)
In inference stage, our GTS models the conditional probability with
$\mathcal{P}(\mathrm{Y}<v|x)=\Phi(\mathbf{h}(x,v)).$ (16)
A.6 Principled guidelines for utilizing GTS
The principled guideline is to align the model assumption (Eq.(11)) as closely as possible with the real data, though this requires manual trial-and-error. Our experience is to first fit the data solely using a mainstream regression model as $\mathbb{Q}$, then analyze the variance, skewness, kurtosis, and outlier distribution of $\mathrm{Y}*\kappa(\mathbb{Q})$ to select a suitable $\kappa$ form, which determines the final GTS model. This significantly reduces training resources when tuning $\kappa$. A more general methodology is left for future work, as stated in the conclusion section.
Appendix B Mathematical Details in Our Approach
B.1 Derivation of transformed MSE’s estimation
The loss function of transformed MSE is exhibited in Eq. (2). Minimizing the loss can result in that
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\big(f(x;\theta)-\mathrm{T}(y)\big)^{2}\Big]$ (17)
$=$ $\min_{\theta}\int_{(x,y)\sim\mathcal{P}\mathrm{(X,Y)}}\big(f(x;\theta)-\mathrm{T}(y)\big)^{2}p(x,y)\mathrm{d}x\mathrm{d}y$
$=$ $\min_{\theta}\int_{x\sim\mathcal{P}(\mathrm{X})}\Big(\int_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big(f(x;\theta)-\mathrm{T}(y)\big)^{2}p(y|x)\mathrm{d}y\Big)p(x)\mathrm{d}x$
$=$ $\min_{\theta}\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[(f(x;\theta)-\mathrm{T}(y))^{2}\big]\Big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\min_{f(x;\theta)}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[(f(x;\theta)-\mathrm{T}(y))^{2}\big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\frac{\partial}{\partial f(x;\theta)}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[(f(x;\theta)-\mathrm{T}(y))^{2}\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[2\times(f(x;\theta)-\mathrm{T}(y))\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,f(x;\theta)=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\big[\mathrm{T}(y)\big].$
Therefore, $f(x;\theta)$ is actually an estimate of $\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\big[\mathrm{T}(y)\big]$.
B.2 Detailed derivation for TranSUN’s estimation
A brief derivation for TranSUN’s estimate is exhibited in Eq. (7, 8). Following Eq. (17), we present detailed derivation for TranSUN’s estimate $\hat{y}|x$ as
$\min_{\theta_{z}}\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}$ (18)
$=$ $\min_{\theta_{z}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\Big]\Big)^{2}\Big]$
$=$ $\min_{\theta_{z}}\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\big]\big)^{2}\big]\Big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\min_{z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\big]\big)^{2}\big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\frac{\partial}{\partial z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\big]\big)^{2}\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[2\times\big(z(x;\theta_{z})-\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\big)\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,z(x;\theta_{z})=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\Big[\frac{y}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}\Big]=\frac{\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]}{|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon}$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\hat{y}|x=z(x;\theta_{z})\cdot\big(|\mathrm{T}^{-1}\big(f(x;\theta)\big)|+\epsilon\big)=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y].$
B.3 Detailed derivation for estimation from the model with S1 scheme
As mentioned in Sec. 3.2 and Table 6, the S1 bias modeling scheme formulates the bias learning loss as
$\mathcal{L}^{\mathrm{T}}_{\mathrm{sun}}(\mathrm{S1})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\Big]\Big)^{2}\Big].$ (19)
Following Eq. (18), we present detailed derivation for its estimate $\hat{y}|x$ as
$\min_{\theta_{z}}\mathcal{L}_{\mathrm{sun}}^{\mathrm{T}}(\mathrm{S1})$ (20)
$=$ $\min_{\theta_{z}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-\mathrm{stop\_grad}\Big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\Big]\Big)^{2}\Big]$
$=$ $\min_{\theta_{z}}\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\big]\big)^{2}\big]\Big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\min_{z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\big]\big)^{2}\big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\frac{\partial}{\partial z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-\mathrm{stop\_grad}\big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\big]\big)^{2}\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[2\times\big(z(x;\theta_{z})-\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\big)\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,z(x;\theta_{z})=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\Big[\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{y}\Big]={\mathrm{T}^{-1}\big(f(x;\theta)\big)}\cdot{\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[\frac{1}{y}]}$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\hat{y}|x=\frac{\mathrm{T}^{-1}\big(f(x;\theta)\big)}{z(x;\theta_{z})}=\frac{1}{\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[\frac{1}{y}]}.$
As shown above, the model prediction $\hat{y}|x$ is actually an estimate of ${1}/{\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[\frac{1}{y}]}$, which is biased.
B.4 Detailed derivation for GTS’s estimation
Owing to the $\mathrm{stop\_grad}$ operation, the optimization of the total loss $\mathcal{L}_{\mathrm{GTS}}$ can be decomposed into two independent parts, as
$\min_{\theta_{q},\theta_{z}}\mathcal{L}_{\mathrm{GTS}}\Leftrightarrow\min_{\theta_{q}}\mathcal{L}_{\mathcal{H}_{q}}+\min_{\theta_{z}}\mathcal{L}_{\mathrm{LTM}}^{\kappa},$ (21)
where $\mathcal{L}_{\mathcal{H}_{q}}$ is the conditional point loss and $\mathcal{L}_{\mathrm{LTM}}^{\kappa}$ is the linear transformation loss. The first term $\min_{\theta_{q}}\mathcal{L}_{\mathcal{H}_{q}}$ learns a conditional point according to Eq. (10), which is independent of $z(x;\theta_{z})$. The second term derives the GTS’s estimate $\hat{y}|x$ via
$\min_{\theta_{z}}\mathcal{L}_{\mathrm{LTM}}^{\kappa}$ (22)
$=$ $\min_{\theta_{z}}\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[\Big(z(x;\theta_{z})-y\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big]\Big)^{2}\Big]$
$=$ $\min_{\theta_{z}}\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-y\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big]\big)^{2}\big]\Big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\min_{z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-y\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big]\big)^{2}\big]$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\frac{\partial}{\partial z(x;\theta_{z})}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[\big(z(x;\theta_{z})-y\cdot\mathrm{stop\_grad}\big[\kappa\big(f(x;\theta_{q})\big)\big]\big)^{2}\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y|}x)}\big[2\times\big(z(x;\theta_{z})-y\cdot\kappa\big(f(x;\theta_{q})\big)\big)\big]=0$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,z(x;\theta_{z})=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}\Big[y\cdot\kappa\big(f(x;\theta_{q})\big)\Big]={\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y]}{\cdot\kappa\big(f(x;\theta_{q})\big)}$
$\Leftrightarrow$ $\forall x\sim\mathcal{P}(\mathrm{X}),\,\hat{y}|x=\frac{z(x;\theta_{z})}{\kappa\big(f(x;\theta_{q})\big)}=\mathbb{E}_{y\sim\mathcal{P}\mathrm{(Y|}x)}[y].$
B.5 Discussion for the theoretical correctness of the unbiasedness metrics (TRE and MRE)
Details are shown in Appendix D.2.2.
B.6 Utilizing GTS for direct probabilistic modeling
Details are shown in Appendix A.5.
Appendix C Challenges of applying external bias correction approaches in industrial recommender systems
Applying external debiasing paradigms from prior works to real-world recommender systems necessitates the development of an additional online post-processing pipeline after the ranking model’s scoring pipeline in the overall online system, which suffers from practical challenges including high development costs, substantial maintenance difficulties, and significant system resource consumption. In addition, since the online calibration pipeline [10] has been commonly utilized in many large-scale recommender systems, an alternative approach is to just let this pipeline itself fit the correction. However, in practice, when the retransformation bias problem becomes severe (e.g. TRE is about 0.6 before calibration), this alternative approach is prone to yield unstable online TRE results after the calibration, which are significantly worse than those of our TranSUN method, with the mean being 180% and the standard deviation being 150% of ours. The instability of calibration results might be caused by the excessively large differences between predicted values and true values, which can even affect the robustness of the calibration pipeline. Therefore, as a practical solution, we propose TranSUN and Generalized TranSUN (GTS), an end-to-end in-model correction method to address the challenges of existing approaches and provide robust model performance of unbiasedness.
Appendix D Details in Experimental Setting
D.1 Detailed setting on synthetic data
D.1.1 Data and metric
We adopt $8$ distributions $\mathcal{H}(\mathrm{Y})$ from three categories: right-skewed, left-skewed, and symmetric. Specifically, we select some classic exponential family instances (e.g. Gamma, Beta, Normal), multimodal distributions (e.g. bimodal uniform), and industry-prevalent zero-inflated distributions (e.g. ZILN [40] with positive rate $p_{+}=0.2$). Detailed implementations are shown in Code 2. For each distribution, we randomly sample $N=10^{6}$ instances. Fig. 2 illustrates the density and some statistics of the sampled $\mathcal{P}_{s}(\mathrm{Y}|x)$. We adopt SRE (Signed Relative Error) as the metric, i.e. ${(\hat{y}-y)}/{y}$ where $\hat{y}$ and $y$ are prediction and ground truth, respectively.
D.1.2 Implementation details
We implement all methods using the Tensorflow toolbox on an M2 Pro CPU with 16 GB. Since the model input $x$ is fixed in $\mathcal{P}_{s}(\mathrm{Y}|x)$, we directly set it to $1.0$. We also implement some advanced classification-based models [23, 38, 21, 9] for comparison, where only their main classification losses are retained. For TPM [23] and CREAD [38], bin number is set to $4$, which is produced via equal-frequency and EAD binning method [38], respectively. MDME [21] adopts an equal-width binning method to determine $2$ sub-distributions, each containing $2$ buckets. WLR [9] is only evaluated on RS-ZIG since it requires substantial zero values as negative samples. We also implement ZILN [40] and OptDist [41], which discard classification branches when not evaluated on RS-ZIG. And OptDist is set to adopt $3$ SDNs. We set batch size to $1024$ and select the optimal learning rate from $\{10^{-i}\}_{1\leq i\leq 3}\cup\{5\times 10^{-i}\}_{1\leq i\leq 4}$. All models are trained for 1 epoch, with each training repeated 10 times with different random seeds to obtain statistical significance, where the $p$-value is calculated by the t-test.
D.2 Detailed setting on real-world data
D.2.1 Datasets
We adopt two public datasets (CIKM16, DTMart) and one exclusive industrial dataset (abbr. Indus). Specifically, CIKM16 aims to predict the dwell time for each session in online searching. We only use the items in the session as input features. The CIKM16 dataset contains 310,302 sessions and 122,991 items. DTMart aims to predict the transaction amount for each order in marketing activities. We use item ID, categories, time taken for procurement, and delivery as input features. The DTMart dataset contains 985,033 orders and 15,546 items. Indus aims to predict the Gross Merchandise Value (GMV) of each order paid in the recommendation scenario of an anonymous e-commerce platform. The Indus dataset contains 132,657,136 orders, 14,538,759 items, and 67,729,869 users. CIKM16 and DTMart datasets are both randomly divided into training and testing samples at a ratio of 4:1, and the Indus dataset is divided at a ratio of 5:1. Interestingly, we found that the Pareto phenomenon occurred in the Indus dataset, where 80% of the GMV was contributed by the top 30% high-priced orders. Consequently, we further conduct evaluation solely on the top high-priced orders in the testing set, abbr. Indus Top test, to better analyze model performance.
D.2.2 Metrics for assessing unbiasedness
As mentioned in Sec. 5.1, we propose Total Ratio Error (TRE) and Mean Ratio Error (MRE) to assess unbiasedness. These metrics are formulated as
$\mathrm{TRE}=\Bigg|\frac{\sum_{i=1}^{M}{\hat{y}_{i}-y_{i}}}{\sum_{i=1}^{M}{\hat{y}_{i}}}\Bigg|,$ (23)
$\mathrm{MRE}=\Bigg|\frac{1}{M}\sum_{i=1}^{M}\frac{{\hat{y}_{i}-y_{i}}}{{\hat{y}_{i}}}\Bigg|,$ (24)
where $\hat{y_{i}}$ is prediction, $y_{i}$} is ground truth, and $M$ is sample number.
Firstly, we theoretically prove that achieving either the optimal TRE or MRE metric is a necessary condition for the model to be unbiased. Specifically, through the definition of the model’s unbiasedness, we can derive that
$\forall x\sim\mathcal{P}(\mathrm{X}),\ \hat{y}(x)=\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}[y]\;\Leftrightarrow\;\forall x\sim\mathcal{P}(\mathrm{X}),\ \mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\hat{y}(x)-y\big]=0,$ (25)
where $\hat{y}(x)=\hat{y}|x$. For the TRE metric, optimizing it means that
$\min\mathrm{TRE}$ (26)
$\Leftrightarrow\Bigg|\frac{\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\hat{y}(x)-y\big]}{\mathbb{E}_{(x,y)\sim\mathcal{D}}[\hat{y}(x)]}\Bigg|=0$
$\Leftrightarrow\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\hat{y}(x)-y\big]=0$
$\Leftrightarrow\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\hat{y}(x)-y\big]\Big]=0$
$\Leftarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\hat{y}(x)-y\big]=0,$
where the second row in Eq. (26) is derived from the definition of TRE using expectation formulation, and the last row comes from Eq. (25). Regarding the MRE metric, optimizing it implies that
$\min\mathrm{MRE}$ (27)
$\Leftrightarrow\bigg|\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[1-\frac{y}{\hat{y}(x)}\Big]\bigg|=0$
$\Leftrightarrow\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\Big[\frac{\hat{y}(x)-y}{\hat{y}(x)}\Big]\Big]=0$
$\Leftrightarrow\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\frac{1}{{\hat{y}(x)}}\cdot\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\Big[{\hat{y}(x)-y}\Big]\Big]=0$
$\Leftarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\hat{y}(x)-y\big]=0,$
where the second row in Eq. (27) is derived from the definition of MRE using expectation formulation, and the last row comes from Eq. (25). Notably, the derivation in Eq. (27) explains the reason that the model prediction $\hat{y}(x)$ is placed in the denominator position of the metric. Specifically, once the MRE is modified to swap the numerator and denominator of each item in it, as
$\mathrm{MRE}^{\mathrm{(wr)}}=\Bigg|\frac{1}{M}\sum_{i=1}^{M}\frac{{y_{i}-\hat{y}_{i}}}{{{y}_{i}}}\Bigg|,$ (28)
then this new metric is no longer suitable for measuring the model unbiasedness, which actually turns to a necessary condition for the estimation of $1/\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\frac{1}{y}\big]$, as
$\min\mathrm{MRE}^{\mathrm{(wr)}}$ (29)
$\Leftrightarrow\bigg|\mathbb{E}_{(x,y)\sim\mathcal{D}}\Big[1-\frac{\hat{y}(x)}{y}\Big]\bigg|=0$
$\Leftrightarrow\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\Big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\Big[1-\frac{\hat{y}(x)}{y}\Big]\Big]=0$
$\Leftrightarrow\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\bigg[{{\hat{y}(x)}}\cdot\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\bigg[{\frac{1}{\hat{y}(x)}-\frac{1}{y}}\bigg]\bigg]=0$
$\Leftarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\bigg[{\frac{1}{\hat{y}(x)}-\frac{1}{y}}\bigg]=0$
$\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \hat{y}(x)=\frac{1}{\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[\frac{1}{y}\big]},$
where the second row in Eq. (29) is derived from the definition of $\mathrm{MRE^{(wr)}}$ using expectation formulation.
Next, we theoretically prove that achieving the optimal NRMSE metric is a necessary and sufficient condition for the model to be unbiased. Specifically, NRMSE is defined as
$\mathrm{NRMSE}=\frac{\sqrt{\sum_{i=1}^{M}(\hat{y}_{i}-y_{i})^{2}}}{\sum_{i=1}^{M}{y_{i}}},$ (30)
and thus optimizing it means that
$\min\mathrm{NRMSE}$ (31)
$\Leftrightarrow\min\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\big(\hat{y}(x)-y\big)^{2}\big]$
$\Leftrightarrow\min\mathbb{E}_{x\sim\mathcal{P}(\mathrm{X})}\big[\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[(\hat{y}(x)-y)^{2}\big]\big]$
$\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \min\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[(\hat{y}(x)-y)^{2}\big]$
$\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \frac{\partial}{\partial\hat{y}(x)}\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[(\hat{y}(x)-y)^{2}\big]=0$
$\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}\big[2\times\big(\hat{y}(x)-y\big)\big]=0$
$\Leftrightarrow\forall x\sim\mathcal{P}(\mathrm{X}),\ \hat{y}(x)=\mathbb{E}_{y\sim\mathcal{P}(\mathrm{Y}|x)}[y],$
where the second row in Eq. (31) is derived from the definition of NRMSE using expectation formulation, and the last row indicates that $\hat{y}(x)$ is an unbiased estimate.
Discussion on not using NRMSE as the primary metric for observing unbiasedness.
Since NRMSE is a necessary and sufficient condition for a model to be unbiased, intuitively it should be preferred over TRE and MRE as the primary metric for observing unbiasedness. However, the NRMSE metric has a significant issue, that its optimal value (i.e. the minimum) is unknown. This problem hinders us from intuitively perceiving the presence and severity of the bias, making it much difficult to verify the capability of different models in model iterations towards eradicating bias. In contrast, although TRE and MRE are both necessary conditions for a model to be unbiased, their optimal values (i.e. the minimum) are known, which is $0.0$. Therefore, they can both clearly indicate whether a model is biased. In our experiments, we take the approach for analyzing unbiasedness where TRE and MRE are initially compared to determine if the model is biased, and then NRMSE is subsequently compared to identify the model with the best unbiasedness. The idea of the analytical approach above is also reflected in our result reporting. For example, NRMSE results are not reported in Table 3 and Fig. 3b3c3d, which is because the gaps among methods in TRE and MRE are both quite pronounced already.
D.2.3 Implementation details
For public datasets, all methods employ an “Embedding+MLP” architecture, where continuous features are discretized, each feature’s embedding dimension is set to $16$, and the MLP has $4$ layers with hidden dimensions of $123$, $64$, and $32$. For Indus, all methods employ a DIN architecture [49], where continuous features are also discretized, and the MLP has $4$ layers with hidden dimensions of $1024$, $512$, and $256$. The bias modeling branch of TranSUN defaults to non-shared parameters on public datasets and shares the embedding layer on Indus. The default setting for $\epsilon$ is $1.0$. For other baselines, WLR [9] folows the implementation in [45]. ZILN [40] and OptDist [41] both discard classification branches. The number of bins in TPM [23] and CREAD [38] is set to $16$. MDME [21] adopts $4$ sub-distributions, each containing $4$ buckets, and OptDist adopts $3$ SDNs. The equal-frequency binning method is adopted for both TPM and MDME, while the EAD method is adopted for CREAD. On public datasets, the SGD optimizer is utilized with a batch size of $512$ and a learning rate of $0.001$. On the Indus, the AdagradDecay optimizer is employed with a batch size of $64$ and a learning rate of $0.01$. All models adopt only one training epoch, and each training is repeated 5 times with different random seeds to obtain statistical significance, where the $p$-value is calculated by the t-test. We implement all methods using the Tensorflow toolbox. The model training on public datasets is performed on an M2 Pro CPU with 16GB, while the training on the Indus dataset consumes resources including 2000 CPUs with 90 GB. In addition, the statistical significance in the results of the online A/B test (Table 9) is calculated by the z-test.
Appendix E More Experimental Results
E.1 Biasedness comparison of TPM with different bin numbers on synthetic data
As shown in Table 2, TPM [23] underestimates the mean on left-skew and overestimates it on right-skew data, likely because it learns the average of quantiles while the bin number is only $4$. The bin number is set so small to amplify the theoretical biasedness problem of TPM, which can be significantly alleviated by increasing the bin number, as shown in Table 14. This conclusion also applies to CREAD [38].
Table 14: Comparison results of TPM with different bin numbers on synthetic data, where $X\#\mathrm{bin}$” indicates the bin number of TPM is $X$. Bold indicates the best results, and gray background indicates that |SRE| <1%.
Models RS-G RS-BU RS-ZIG LS-B LS-BU SM-U SM-TN SM-BU
[…]
E.2 Details for online A/B test
In Taobao App (DAU > 300M), our LogSUN model is deployed as GMV prediction models in both product and short video recommendation scenarios in Guess What You Like domain in the homepage, which serve as pointwise ranking models for GMV target in the ranking stage of the recommender systems to optimize object distribution. Notably, in the short video recommendation scenario on the e-commerce platform, all the short videos are product description videos, each of which links to a certain product, therefore, the GMV of a short video actually refers to the transaction value of the product described in the video after the short video is clicked.
In the A/B test, our LogSUN is compared against the previous online base (LogMAE) and the commonly used LogMSE. Two common metrics, GMV and GMVPI (GMV Per Impression), are adopted to assess online performance. Due to the large variability of these GMV-related metrics, we allocated either substantial traffic or a sufficiently long period for A/B tests to obtain statistically significant results. Specifically, we conduct two A/B tests for one week in the product recommendation scenario, with LogSUN serving 15% of the total traffic, LogMAE 20%, and LogMSE 9%. And we conduct one A/B test for one month in the short video recommendation scenario, with LogSUN serving 6% of the total traffic, and LogMAE 6%.
As shown in Table 9, LogSUN achieves significant gains (e.g. > +1.0% GMV gains) over the baselines in both recommendation scenarios. We attribute the consistent gains in online metrics to the following two reasons. The first reason is that LogSUN significantly improves the predictive accuracy of samples in each target bin by eradicating the retransformation bias, as shown in Figure 3c. The second is that LogSUN empirically improves the ranking accuracy of the top high-priced samples (e.g. the improvements in AUC on Indus Top test, as shown in Table 8), which contribute 80% of the total GMV though they are small in number. On the other hand, we observed that the GMV gain in the short video recommendation scenario was much larger than that of the product. This is possibly because the retransformation bias problem is much more severe in the short video recommendation scenario (TRE is about 0.6352, nearly 140% of the product’s). Now, LogSUN has been deployed in the two core recommendation scenarios to serve the major traffic.
E.3 Comparison of GTS instance models on synthetic data
Some common $\mathcal{L}_{\mathcal{H}_{q}}$ and $\kappa$ are selected to devise different instances of GTS, providing reference results for its application. As exhibited in Table 15, in terms of $\mathcal{L}_{\mathcal{H}_{q}}$, MSE loss is the most versatile loss as it fastly converges to the unbiased solution in all distributions, whose $\mathbb{Q}$ corresponds to the arithmetic mean point. Among other losses, MAE presents better convergence on right-skewed data, while MSPE and MAPE both converge more effectively on left-skewed and symmetric data. Regarding function $\kappa$, all candidates exhibit unbiasedness consistent with TranSUN. The performance of $\kappa$ on real-world data is shown in Table 10.
Table 15: Comparison results of different GTS instances ($\mathrm{T}(y)=\ln(y+1)$) on synthetic data . Red boldface indicates that |SRE| >1%.
$\mathcal{L}_{\mathcal{H}_{q}}$ $\kappa(\cdot)$ RS-G RS-BU RS-ZIG LS-B LS-BU SM-U SM-TN SM-BU
MSE $\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.0014 0.0013 -0.0010 0.0014 0.0033 -0.0038 -0.0017 -0.0058
MAE $\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.0042 0.0010 -0.0088 -0.0012 0.0242 0.0177 0.0083 0.0010
MSPE $\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.0083 0.0340 -0.0048 -0.0064 -0.0044 -0.0074 0.0059 0.0023
MAPE $\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.0103 0.0191 -0.0018 -0.0087 -0.0016 -0.0084 0.0047 -0.0061
MSE $\kappa(u)=1/(|u|+\epsilon)$ 0.0055 0.0044 -0.0011 -0.0020 -0.0032 -0.0034 0.0036 0.0041
MSE $\kappa(u)=|u|$ 0.0018 0.0024 -0.0057 -0.0014 -0.0018 -0.0028 -0.0012 -0.0044
E.4 Architectural sensitivity on real-world data
Since our method introduces an extra bias learning branch to transformed MSE, we explore the impact of different parameter sharing schemes between the added branch and the original model, which is conducted on the LogSUN model. In Table 16, sharing parameters from bottom layers (e.g. embedding layers) has minimal impact on performance, whereas further sharing more parameters within MLP results in a noticeable performance degradation, which might be inferred that the added task primarily affects the downstream MLP-modeled fitting function rather than the distribution of latent embedding.
[…]
Parameter sharing schemes TRE $\downarrow$ MRE $\downarrow$ NRMSE $\downarrow$ NMAE $\downarrow$ XAUC $\uparrow$
[…]
$1/1$ Embedding + $2/3$ MLP 0.0490 0.0686 0.5684 0.4455 0.6702
$1/1$ Embedding + $3/3$ MLP 0.1063 0.0955 0.5951 0.4775 0.6658
E.5 Comparison of GTS instances with different $\kappa$ functions on real-world data
As shown in Table 10, compared to the one used in TranSUN, other $\kappa$ candidates all bring about poorer loss convergence, which is probably because the variances of their $\kappa\cdot y$ are all too large (Table 17), leading to an excessively steep optimization landscape for $\theta_{z}$.
Table 17: Comparison of the variances of $\kappa\cdot y$ in GTS instances with different $\kappa$ functions ($\mathrm{T}(y)=\ln(y+1)$ and $\mathcal{L}_{\mathcal{H}_{q}}$ is set to MSE loss) on CIKM16 train.
$\kappa(\cdot)$ $\mathrm{Var}(\kappa\cdot y)$ $\mathrm{Var}(\kappa\cdot y)/(\mathbb{E}[\kappa\cdot y])^{2}$
$\kappa(u)=1/(|\mathrm{T}^{-1}(u)|+\epsilon)$ 0.6329 0.3610
$\kappa(u)=1/(|u|+\epsilon)$ 417.67 0.3829
$\kappa(u)=|u|$ 461790.3 0.4447
Appendix F Broader Impacts
We think that there is no societal impact of the work performed. Though our method has been applied in two core recommendation scenarios on a leading e-commerce platform (DAU > 300M) to serve the major online traffic, there are no obvious paths that lead to potential harm, since our work just presents an unbiased regression methodology and is not tied to particular applications.
Code 1: TensorFlow-style implementation for our TranSUN and GTS.
1import numpy as np
2import tensorflow.compat.v1 as tf
3
4tf.disable_v2_behavior()
5from tensorflow.keras.layers import Dropout
6
7
8def transun_model(x, y_label, mlp_units, trans,
9 dropout, feat_num, feat_dim, rmask, imask, reuse, is_training,
10 **kwargs):
11 with tf.variable_scope(’transun’, reuse=reuse):
12 # embed & pool
13 embedding_param = tf.get_variable(name="embtable", shape=[feat_num, feat_dim])
14 feas = tf.nn.embedding_lookup(ids=x, params=embedding_param) # (bsz,20,feat_dim)
15 requests = tf.reduce_sum(feas[:, :10, :] * tf.expand_dims(rmask, -1), 1) / tf.reduce_sum(rmask, 1, keepdims=True) # avg pooling
16 items = tf.reduce_sum(feas[:, 10:, :] * tf.expand_dims(imask, -1), 1) / tf.reduce_sum(imask, 1, keepdims=True)
17 feas = tf.concat([requests, items], 1) # (bsz,2*feat_dim)
18 # mlp, main branch
19 drpt = Dropout(rate=dropout)
20 fc = feas
21 for idx, num_unit in enumerate(mlp_units):
22 fc = linear(fc, num_unit, f"dense{idx}", reuse)
23 fc = tf.nn.relu(fc)
24 fc = drpt(fc, training=is_training)
25 logits = linear(fc, 1, "output", reuse)
26 # mlp, sun branch
27 fc1 = feas
28 for idx, num_unit in enumerate(mlp_units):
29 fc1 = linear(fc1, num_unit, f"sun{idx}", reuse)
30 fc1 = tf.nn.relu(fc1)
31 fc1 = drpt(fc1, training=is_training)
32 sun = linear(fc1, 1, "sun_out", reuse)
33
34 bias = 1.0
35
36 if trans == ’log’:
37 ori_pred = tf.math.exp(tf.math.minimum(10.0,tf.nn.relu(logits))) - 1
38 final_pred = (ori_pred + bias) * tf.nn.relu(sun)
39 elif trans == ’sqrt’:
40 ori_pred = tf.nn.relu(logits) ** 2
41 final_pred = (ori_pred + bias) * tf.nn.relu(sun)
42 elif trans == ’atan’:
43 ori_pred = tf.math.tan(tf.clip_by_value(logits, 0.0, np.atan(10000)))
44 final_pred = (ori_pred + bias) * tf.nn.relu(sun)
45 elif trans == ’tanh’:
46 ori_pred = tf.math.atanh(tf.clip_by_value(logits, 0.0, np.tanh(800)))
47 final_pred = (ori_pred + bias) * tf.nn.relu(sun)
48 else:
49 raise NotImplementedError
50
51 FUSE_THRES = 54.0
52 fused_pred = tf.where(ori_pred < FUSE_THRES, ori_pred, final_pred)
53
54 if is_training:
55 if trans == ’log’:
56 sun_label=tf.stop_gradient(y_label / (tf.exp(logits) - 1 + bias))
57 mse_loss = tf.reduce_sum(tf.square(tf.log(y_label + 1) - logits))
58 sun_loss = tf.reduce_sum(tf.square(sun - sun_label))
59 elif trans == ’sqrt’:
60 sun_label = tf.stop_gradient(y_label / (logits ** 2 + bias))
61 mse_loss = tf.reduce_sum(tf.square(tf.math.sqrt(y_label) - logits))
62 sun_loss = tf.reduce_sum(tf.square(sun - sun_label))
63 elif trans == ’atan’:
64 sun_label=tf.stop_gradient(
65 y_label / (tf.math.tan(tf.clip_by_value(logits, 0.0, np.atan(10000))) + bias))
66 mse_loss = tf.reduce_sum(tf.square(tf.math.atan(y_label) - logits))
67 sun_loss = tf.reduce_sum(tf.square(sun - sun_label))
68 elif trans == ’tanh’:
69 sun_label = tf.stop_gradient(y_label / (tf.math.atanh(tf.clip_by_value(logits, 0.0, np.tanh(800))) + bias))
70 mse_loss = tf.reduce_sum(tf.square(tf.math.tanh(y_label) - logits))
71 sun_loss = tf.reduce_sum(tf.square(
72 sun - sun_label)
73 )
74 else:
75 raise NotImplementedError
76
77 rank_loss0 = _rank_loss(sun,sun_label)
78 rank_loss1 = _rank_loss(logits, y_label)
79 rank_loss = rank_loss0 #+ rank_loss1
80 return {’loss’: mse_loss
81 + sun_loss
82 + rank_loss,
83 ’mse_loss’: mse_loss,
84 ’sun_loss’: sun_loss,
85 # ’rank_loss’:rank_loss,
86 ’gt_mean’: tf.reduce_mean(y_label),
87 ’sun_gt_mean’: tf.reduce_mean(sun_label),
88 ’sun_pred_mean’: tf.reduce_mean(sun),
89 ’pred_mean’: tf.reduce_mean(final_pred),
90 ’ori_pred_mean’: tf.reduce_mean(ori_pred),
91
92 ’pred’: final_pred,
93 ’gt’: y_label,
94 ’ori_pred’: ori_pred
95 }
96 else:
97 return {"pred": final_pred,
98 "fused_pred": fused_pred,
99 "ori_pred": ori_pred}
100
101
102def gts_model(x, y_label, mlp_units,
103 dropout, feat_num, feat_dim, rmask, imask, reuse, is_training, location=’mse’, k_func=0, trans=’log’, **kwargs):
104 with tf.variable_scope(’gts’, reuse=reuse):
105 embedding_param = tf.get_variable(name="embtable", shape=[feat_num, feat_dim])
106 feas = tf.nn.embedding_lookup(ids=x, params=embedding_param) # (bsz,20,feat_dim)
107 requests = tf.reduce_sum(feas[:, :10, :] * tf.expand_dims(rmask, -1), 1) / tf.reduce_sum(rmask, 1, keepdims=True) # avg pooling
108 items = tf.reduce_sum(feas[:, 10:, :] * tf.expand_dims(imask, -1), 1) / tf.reduce_sum(imask, 1, keepdims=True)
109 feas = tf.concat([requests, items], 1) # (bsz,2*feat_dim)
110 # mlp, main branch
111 drpt = Dropout(rate=dropout)
112 fc = feas
113 for idx, num_unit in enumerate(mlp_units):
114 fc = linear(fc, num_unit, f"dense{idx}", reuse)
115 fc = tf.nn.relu(fc)
116 fc = drpt(fc, training=is_training)
117 logits = linear(fc, 1, "output", reuse)
118 fc = feas
119 for idx, num_unit in enumerate(mlp_units):
120 fc = linear(fc, num_unit, f"dense_sun{idx}", reuse)
121 fc = tf.nn.relu(fc)
122 fc = drpt(fc, training=is_training)
123 sun = linear(fc, 1, "sun", reuse)
124
125 assert location in [’mse’, ’mae’, ’mspe’, ’mape’]
126 assert k_func in [0, 1, 2]
127 assert trans in [’log’]
128
129 if trans == ’log’:
130 label = tf.log(y_label + 1)
131 ori_pred = tf.exp(logits) - 1
132
133 if location == ’mse’:
134 location_loss = tf.reduce_mean(tf.square(logits - label))
135 elif location == ’mae’:
136 location_loss = tf.reduce_mean(tf.abs(logits - label))
137 elif location == ’mspe’:
138 location_loss = tf.reduce_mean(tf.square((logits - label) / (label + 1)))
139 elif location == ’mape’:
140 location_loss = tf.reduce_mean(tf.abs((logits - label) / (label + 1)))
141
142 if k_func == 0:
143 sun_label = tf.stop_gradient(y_label * tf.exp(-logits))
144 y_pred = tf.exp(tf.math.minimum(10.0,logits)) * tf.nn.relu(sun)
145 elif k_func == 1:
146 sun_label = tf.stop_gradient(y_label / (1 + tf.abs(logits)))
147 y_pred = (1 + tf.abs(logits)) * tf.nn.relu(sun)
148 elif k_func == 2:
149 sun_label = tf.stop_gradient(y_label * tf.abs(logits))
150 y_pred = tf.nn.relu(sun) / (1e-12+tf.abs(logits))
151 sun_loss = tf.reduce_mean(tf.square(sun - sun_label))
152
153 loss = location_loss + sun_loss
154 if is_training:
155 return {’loss’: loss,
156 # +rank_loss,
157 ’location_loss’: location_loss,
158 ’sun_loss’: sun_loss,
159 # ’rank_loss’:rank_loss,
160 ’gt_mean’: tf.reduce_mean(y_label),
161 ’sun_gt_mean’: tf.reduce_mean(sun_label),
162 ’sun_pred_mean’: tf.reduce_mean(sun),
163 ’pred_mean’: tf.reduce_mean(y_pred),
164 ’ori_pred_mean’: tf.reduce_mean(ori_pred),
165 # ’grad_main_emb’: tf.gradients([sun_loss],embedding_param),
166 # ’grad_sun_emb’: tf.gradients([sun_loss],embedding_param),
167 # ’grad_main_logit’: tf.gradients([sun_loss],logits),
168
169 ’pred’: y_pred,
170 ’gt’: y_label,
171 ’ori_pred’: ori_pred
172 }
173 else:
174 return {"pred": y_pred, "ori_pred": ori_pred}
175
176
177def linear(inputs, units, name, reuse=None):
178 with tf.variable_scope(name, reuse=reuse):
179 input_dim = inputs.get_shape()[-1].value
180 weights = tf.get_variable(
181 ’weights’, shape=[input_dim, units], initializer=he_initializer)
182 biases = tf.get_variable(
183 ’biases’, shape=[units], initializer=tf.zeros_initializer())
184 outputs = tf.matmul(inputs, weights) + biases
185 return outputs
186
187
188def _rank_loss(logits, labels):
189 """
190␣␣␣␣@params:
191␣␣␣␣␣␣␣␣-logits:(bsz,1)
192␣␣␣␣␣␣␣␣-labels:(bsz,1)
193␣␣␣␣"""
194 pairwise_logits = logits - tf.transpose(logits)
195 pairwise_mask = tf.greater(labels - tf.transpose(labels), 0)
196 pairwise_logits = tf.boolean_mask(pairwise_logits, pairwise_mask)
197 pairwise_psudo_labels = tf.ones_like(pairwise_logits)
198 rank_loss = tf.reduce_mean(tf.nn.sigmoid_cross_entropy_with_logits(
199 logits=pairwise_logits,
200 labels=pairwise_psudo_labels
201 ))
202 # set rank loss to zero if a batch has no positive sample.
203 rank_loss = tf.where(tf.is_nan(rank_loss), tf.zeros_like(rank_loss), rank_loss)
204 return rank_loss
Code 2: Python codes for generating synthetic data.
1import numpy as np
2from scipy.stats import beta, gamma, uniform, skew, truncnorm
3
4n_samples = 1000000
5zero_inflate_ratio = 0.8
6
7def generate_distributions():
8 distributions = [] # [(title, samples, mean, skewn, var, caption)]
9
10 # (a) RS-G
11 d = gamma(a=2, scale=1)
12 samples = d.rvs(n_samples)
13 true_mean = d.mean()
14 skewn = d.stats(moments=’s’)
15 var = d.var()
16 distributions.append(("Right-Skewed:Gamma", samples, true_mean, skewn, var, ’(a)RS-G’))
17
18 # (b) RS-BU
19 low_samples = uniform(loc=1, scale=10).rvs(int(n_samples * 0.9))
20 high_samples = uniform(loc=90, scale=10).rvs(int(n_samples * 0.1))
21 samples = np.concatenate([low_samples, high_samples])
22 true_mean = 0.9 * 6 + 0.1 * 95
23 skewn = float(skew(samples))
24 var = np.var(samples)
25 distributions.append(("Right-Skewed:Bi-Uniform", samples, true_mean, skewn, var,’(b)RS-BU’))
26
27 # (c) RS-ZIG
28 base_samples = gamma(a=2, scale=1).rvs(n_samples)
29 zero_mask = np.random.rand(n_samples) < zero_inflate_ratio
30 samples = base_samples * (1 - zero_mask)
31 true_mean = gamma(a=2, scale=1).mean() * (1 - zero_inflate_ratio)
32 skewn = skew(samples)
33 var = np.var(samples)
34 distributions.append(("Right-Skewed:Zero-InflatedGamma", samples, true_mean, skewn, var,’(c)RS-ZIG’))
35
36 # (d) LS-B
37 d = beta(a=3, b=1.5)
38 samples = d.rvs(n_samples)
39 true_mean = d.mean()
40 skewn = d.stats(moments=’s’)
41 var = d.var()
42 distributions.append(("Left-Skewed:Beta", samples, true_mean, skewn, var,’(d)LS-B’))
43
44 # (e) LS-BU
45 low_samples = uniform(loc=1, scale=10).rvs(int(n_samples * 0.1))
46 high_samples = uniform(loc=90, scale=10).rvs(int(n_samples * 0.9))
47 samples = np.concatenate([low_samples, high_samples])
48 true_mean = 0.1 * 6 + 0.9 * 95
49 skewn = float(skew(samples))
50 var = np.var(samples)
51 distributions.append(("Left-Skewed:Bi-Uniform", samples, true_mean, skewn, var,’(e)LS-BU’))
52
53 # (f) SM-U
54 d = uniform(loc=0, scale=100)
55 samples = d.rvs(n_samples)
56 true_mean = d.mean()
57 skewn = d.stats(moments=’s’)
58 var = d.var()
59 distributions.append(("Symmetric:Uniform", samples, true_mean, skewn, var,’(f)SM-U’))
60
61 # (g) SM-TN
62 a, b = (0 - 50) / 10, (100 - 50) / 10
63 samples = truncnorm(a, b, loc=50, scale=10).rvs(n_samples)
64 true_mean = 50
65 skewn = skew(samples)
66 var = d.var()
67 distributions.append(("Symmetric:Trunc-Normal", samples, true_mean, skewn, var,’(g)SM-TN’))
68
69 # (h) SM-BU
70 low_samples = uniform(loc=1, scale=10).rvs(int(n_samples * 0.5))
71 high_samples = uniform(loc=90, scale=10).rvs(int(n_samples * 0.5))
72 samples = np.concatenate([low_samples, high_samples])
73 true_mean = 0.5 * 6 + 0.5 * 95
74 skewn = float(skew(samples))
75 var = np.var(samples)
76 distributions.append(("Symmetric:Bi-Uniform", samples, true_mean, skewn, var,’(h)SM-BU’))
77
78 return distributions
79
80
81dist = generate_distributions()